OpenStudy (anonymous):

On a backpacking trip through Europe, Ethan bought snacks for 10 Swiss francs. If the foreign exchange rate between the U.S. dollar and Swiss franc is 1:2, how much did Ethan spend in dollars? $5

OpenStudy (texaschic101):

1/2 = x/10 2x = 10 x = 5 your correct